I don't know how to get the answer. Don't know where to start. If you can help it would be greatly appreciated. I.m a 62 yr old grandmom and sometimes my mind can't comprehend like I wanted it to.
State the domain of f(x)= x-3/x square -16

Found 2 solutions by venugopalramana, uma:
Answer by venugopalramana(3286)

I don't know how to get the answer. Don't know where to start. If you can help it would be greatly appreciated. I.m a 62 yr old grandmom and sometimes my mind can't comprehend like I wanted it to.
OH!ARE YOU PREPARING FOR SOME COURSE OR HELPING YOUR WARDS.EITHER WAY IT IS GREAT TRY WORTH APPLAUDING.LET ME EXPLAIN.
WE CLL X AS INDEPENDENT VARIABLE AND F(X) OR Y IT IS OFTEN CALLED AS DEPENDENT VARIABLE ON X.THE RELATION BETWEEN THE TWO IS THE GIVEN EQUATION
Y = F(X) = (X-3)/(X SQUARE-16)
BY DOMAIN ,WE MEAN THE VALUES X CAN TAKE TO MAKE THE RELATION MEANING FULL.
HERE THERE IS DENOMINATOR IN THE FUNCTION .IF IT BECOMES ZERO THE FUNCTION BECOMES INDETERMINABLE.SO WE SAY X CAN TAKE ANY VALUE OTHER THAN THAT WHICH MAKES THE DENOMINATOR ZERO..THAT IS X SQUARE -16 =0
OR...X = +4 ...AND -4
SO THE DOMAIN IS ALL REAL VALUES EXCEPT X=+4 AND -4.
IT IS REPRESENTED BY THE FOLLOWING NOTATION
(-INFINITY,-4)U (-4,+4)U(+4,INFINITY)
HOPE IT IS CLEAR
State the domain of f(x)= x-3/x square -16

Answer by uma(370)

f(x) = (x-3)/(x^2-16)
To find the domain we need to see which all values the function cannot take.
The given function is not defined when x^2 - 16 = 0
==> x^2 = 16
==> x = +4, -4
So domain = allreal numbers except 4, - 4.
==> D = R - {4, -4}

Good Luck!!!